A certain constant volume gas thermometer recalls pressure 1000mmHg anddd 1366mmHg at the ice and steam point respectively.The pressure at temperature (T) is PT.Show that the Celsius scale temperature is given as T=(PT÷1000 -1)
1
Expert's answer
2019-07-17T09:51:59-0400

Gay-Lussac's law is given by formula

"\frac{p_1}{T_1}=\frac{p_2}{T_2}(1)"

In our case, p1=1000 mm, p2=1000 mm,

Using }(1) we get

T=(PT÷1000 -1)